OpenStudy (anonymous):

How would I solve this problem? log2^x + log2^5 = 4

OpenStudy (whpalmer4):

Adding the logarithms of two quantities is equivalent to taking the logarithm of their product. \[\log a + \log b = \log (a*b)\]

OpenStudy (whpalmer4):

Do we know the base of the log here? I'm wondering if this is really \[\log 2^x\] or maybe \[\log_{2}x\]

OpenStudy (anonymous):

It's the second option.

OpenStudy (whpalmer4):

ah, yes, that makes a big difference :-)

OpenStudy (whpalmer4):

My first statement still holds, so we would have \[\log_{2}x + \log_{2}5 = 4\]\[\log_{2}(5x) = 4\]Raise the log base to each side to eliminate the log and we haev \[2^{\log_{2}{5x}} = 2^4\]\[5x=2^4\]

OpenStudy (whpalmer4):

that ^ in the problem statement threw us off! :-)

OpenStudy (whpalmer4):

if you had written log2 x + log2 5 = 4 I would have understood you meant the log base 2, but this looked like you really meant 2^x and 2^5. not too hard to solve the other version (assume base 10 for the log) \[\log 2^x + \log 2^5 = 4\]\[\log {2^{x+5}} = 4\]Raise base to both sides \[2^{x+5} = 10^4\]Now take log base 2 of both sides \[x+5 = \log_{2}10^4\]\[x = \log_{2}10^4 - 5 \approx 8.28771\]

OpenStudy (anonymous):

Thank you. c: